MASS-ANALYZED THRESHOLD IONIZATION AND STRUCTURES OF M_{3}C_{2}(M=Sc, La)

\underline{LU WU}, ROUDJANE MOURAD and D. S. YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

M_{3}C_{2} (M=Sc, La) clusters are produced by laser vaporization in a pulsed metal-cluster source and identified by photoionization mass spectrometry. Vibrationally resolved ion spectra are obtained with mass-analyzed threshold ionization (MATI) spectroscopy. The MATI spectra of M_{3}C_{2} (M=Sc, La) exhibit a weak 0-0 transition, indicating a significant geometry difference between the neutral and ionized clusters. The ionization energies of Sc_{2}C_{2} and La_{3}C_{2} are measured to be 36398 (5) and 30051(5) cm^{−1}, respectively. In addition, the spectra of the two clusters display a number of vibrational intervals that are associated with M_{3} deformations. Preliminary data analysis shows that both clusters have a C_{2v} bi-pyramid structure in the neutral state and a D_{3h} bi-pyramid structure in the ion state, and the spectra may be assigned to the ^{1}A′_{1} (D_{3h})← ^{2}B_{2} (C_{2v}) transitions.